OpenStudy (anonymous):

help please! if f(x) =x^2 +6 find and simplify the following: a) f(t+5) b) f(t^2+5) c f(2) d) 2f(t) e (f(t))^2 +5

OpenStudy (anonymous):

do you need the answer or the work?

OpenStudy (anonymous):

a. t^2+10t+31 b. t^4+10t+31 c. 10 d. 2t^2+12 e. t^4+12t^2+41

OpenStudy (anonymous):

the work too i want to understand

OpenStudy (anonymous):

first plug n the value of x into every place where you see an x. okay for every question factor like this: a. F(t+5)=x^2+6 =(t+5)^2+6 =(t^2+10t+25)+6 =t^2+10t+25+6 =t^2+10t+31